A valószínűségszámítás II. gyakorlaton való részvétel

dátum:  IX. 9. IX. 16. IX. 23. IX. 30. X. 7. X. 14. X.21 X.28

diákok:

Barják Tamás  +  +  +  +  +  −  −  +
Barkován Anett  +  +  +  +  +  +  +  −
Bihari Szabina  +  −  +  +  +  +  +  +
Bócsi Judit  +  +  −  +  −  −  −  −
Bodnár Zsuzsanna  +  +  +  −  +  +  +  +
Breszkócs Péter Lajos  +  +  +  +  −  +  +  +
Flaskár Ibolya  +  +  +  −  −  +  −  −
Gajdics László  +  +  +  +  +  +  +  +
Herczik Bettina  −  −  +  +  +  +  −  +
Hegedűs Judit  −  +  +  +  +  +  +  −
Husiová Anikó  +  +  +  +  +  +  +  −
Karakas Éva  +  +  +  +  +  +  +  +
Kiss Annamária  −  −  +  −  +  −  −  −
Kovács Szilvia  +  +  +  +  +  +  +  +
Kovács Viktor  +  +  +  +  +  +  +  +
Kovács Zoltán  +  +  −  +  −  +  +  −
Kozma Nikoletta  −  +  +  +  +  +  −  −
Kövér Ildikó  +  +  +  +  +  +  +  +
Kurtos Adrienn  +  −  +  +  +  +  +  −
Milkovszki Tamás  +  +  −  +  +  +  −  +
Molnár Péter Miklós  −  +  +  +  +  +  −  −
Péter Antal  +  −  −  −  +  −  −  −
Sárközi Borbála  +  +  +  +  +  +  +  +
Sebestény Csilla  +  +  +  +  +  −  +  −
Tóth Andrea  +  −  +  +  +  +  +  +
Tóth Eszter  +  −  +  +  +  −  +  +
Tóth István Márk  +  +  +  +  −  +  +  −
Vigh Csilla  +  +  +  +  −  −  +  +

dátum:  XI. 04. XI. 11. XI. 18. XI. 25. XII. 2. XII. 9. XII. 16.

diákok: Barják Tamás Barkován Anett Bihari Szabina Bócsi Judit Bodnár Zsuzsanna Breszkócs Péter Lajos Flaskár Ibolya Gajdics László Herczik Bettina Hegedűs Judit Husiová Anikó Karakas Éva Kiss Annamária Kovács Szilvia Kovács Viktor Kovács Zoltán Kozma Nikoletta Kövér Ildikó Kurtos Adrienn Milkovszki Tamás Molnár Péter Miklós Péter Antal Sárközi Borbála Sebestény Csilla Tóth Andrea Tóth Eszter Tóth István Márk Vigh Csilla